You are conducting a study to see if the accuracy rate for fingerprint identification is significantly more than 0.66. You use a significance level of α=0.001α=0.001.

      H0:p=0.66H0:p=0.66
      H1:p>0.66H1:p>0.66

You obtain a sample of size n=434n=434 in which there are 297 successes.

What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic =

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value =

The p-value is...

  • less than (or equal to) αα
  • greater than αα



This test statistic leads to a decision to...

  • reject the null
  • accept the null
  • fail to reject the null



As such, the final conclusion is that...

  • There is sufficient evidence to warrant rejection of the claim that the accuracy rate for fingerprint identification is more than 0.66.
  • There is not sufficient evidence to warrant rejection of the claim that the accuracy rate for fingerprint identification is more than 0.66.
  • The sample data support the claim that the accuracy rate for fingerprint identification is more than 0.66.
  • There is not sufficient sample evidence to support the claim that the accuracy rate for fingerprint identification is more than 0.66.You are conducting a study to see if the accuracy rate for fingerprint identification is significantly more than 0.66. You us
Answer #1

Solution :

Given that,

p_{0} = 0.66

1 - p_{0} = 0.34

n = 434

x = 297

Level of significance = \alpha = 0.001

Point estimate = sample proportion = \hat p = x / n = 0.684

This a right (One) tailed test.

Ho: p = 0.66

Ha: p > 0.66

Test statistics

z = (\hat p - p_{0} ) / \sqrt{} p_{0}*(1-p_{0}) / n

= ( 0.684 - 0.66) / \sqrt{} (0.66*0.34) / 434

= 1.055

P-value = P(Z>z)

= 1 - P(Z <z )

= 1- P(Z < 1.055)

= 1 - 0.8543

= 0.1447

The p-value is p = 0.1447, and since p = 0.1447 > 0.001, it is concluded that the null hypothesis is fail to reject.

The p-value is greater than \alpha .

Fail to reject the null hypothesis.